A heavy machine in the factory produces a sound intensity level of 130 dB in relation to the audibility threshold. Determine the level of sound intensity that will be produced by two such machines that would work simultaneously.


1
Expert's answer
2021-03-26T11:41:03-0400

The sound intensity level of a single machine is


"\beta_1=10\log_{10}\frac{I}{I_0}."

In this equation, I0 is the audibility threshold equal to 10-12 W/m2.

So, for a single machine,


"I=I_0·10^{\beta_1/10}."

For two similar machines working simultaneously:


"\beta_2=10\log_{10}\frac{2I}{I_0}=10\log_{10}\frac{2I_0·10^{\beta_1/10}}{I_0}=\\\space\\
=10\log_{10}[2·10^{\beta_1/10}]=133\text{ dB}."